OpenStudy (anonymous):

Solve for x: 8^x = 1/6root(2)

OpenStudy (anonymous):

I'm guessing I should change the RHS to (1/2)^(1/6.) Am I right?

OpenStudy (anonymous):

lol tkhunny saves the day again!

OpenStudy (anonymous):

Hello to you as well, sasogeek

OpenStudy (sasogeek):

hi :)

OpenStudy (tkhunny):

First, I'm not clear on your notation. Is it \(8^{x} = \dfrac{1}{6}\sqrt{2}\) or maybe \(8^{x} = \sqrt[6]{2}\) or something else?

OpenStudy (anonymous):

The latter.

OpenStudy (anonymous):

Sorry. I don't know how to type expressions like this into the field.

OpenStudy (anonymous):

Anyway..I imagine I'm supposed to end up with 2^3^x = (1/2)^(1/6)

OpenStudy (anonymous):

And then 2^3x = 2^(-1/6)

OpenStudy (tkhunny):

No worries. Learning just a little LaTeX can go a long way. You just have to find other ways to communicate without it. Click the [Equation] button, below, and do some experimenting.

OpenStudy (anonymous):

Alright, will do :). So...is the answer just (-1/6)/2?

OpenStudy (tkhunny):

Excellent! 2^(3x) is the real key to this problem's simplest solution. Same bases! If they are to be equal, the exponents must be equal. 3x = -1/6 if we started with \(\dfrac{1}{\sqrt[6]{2}}\) or 3x = 1/6 if we started with \(\sqrt[6]{2}\) or

OpenStudy (anonymous):

Oh...of course. So it's (-1/6)/3, or -1/18

OpenStudy (anonymous):

haha Thank you very much for the medal sasogeek, however undeserved it may be!

OpenStudy (anonymous):

Thank you for all of your help!

OpenStudy (sasogeek):

it is well deserved :) you worked your way through to get the answer ;) and u got it right as well :)

OpenStudy (anonymous):

Well, actually I got it wrong and then tkhunny guided me towards the correct answer, but I'm glad you think so!

OpenStudy (anonymous):

A good night to all!

OpenStudy (anonymous):

Though I'm sure I'll be back.

OpenStudy (sasogeek):

sure he did :) and that was great of him/her lol but at the end of the day, i think you deserve the medal for your effort :)

OpenStudy (anonymous):

Well, once again, thank you very much!